We saw a strange striation pattern in some images it looked like the old syncing problem, but really was a problem with the controllers. We resolved the problem by cycling the power of the Monsoon controllers.
The controller problem returned starting at image obj_0478. WE moved the telescope to zenith and I went to the Cass cage to power cycle the Monsoon electronics.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2023-09-29 | 8.5 | 0.0 | 1.5 | 0.0 | 0.0 | 10 |
Links
[1] mailto:sean.points@noirlab.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42930/delete?destination=print/content/enr-ctio-blanco-4-m-2023-09-29
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42930/edit?destination=print/content/enr-ctio-blanco-4-m-2023-09-29
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/20928?destination=print/content/enr-ctio-blanco-4-m-2023-09-29
Had steady focus for he first few hours of the night and the delivered image quality was 0.8¨ (2 pix). Used that time to calculate the focus offsets among the different filters. Then performed tests to characterize the narrow band filters (164, 2124, and 2168).